OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

Is it (2 square root of 2, 135°), (-2 square root of 2, 315°)?

OpenStudy (anonymous):

[let x=rcos theta ,y =r sin theta ] \[2=r \cos \theta ,-2=r \sin \theta \] square and add \[4+4=r ^{2}\left( \cos ^{2}\theta+\sin ^{2}\theta \right),r ^{2}=8,r=2\sqrt{2}\] divide \[\tan \theta =-1=-\tan \frac{ \pi }{4 }=\tan \left( \pi-\frac{ \pi }{4 } \right),\tan \left( 2\pi-\frac{ \pi }{4 } \right)\] \[\theta=\frac{ 3\pi }{4 }=135 and \theta=\frac{ 7\pi }{4 }=315\] polar co-ordinates are \[(2\sqrt{2},135) and (2\sqrt{2},315)\]

OpenStudy (anonymous):

r is always positive.

OpenStudy (anonymous):

Well it says one has to be negative.

OpenStudy (anonymous):

Thanks btw

OpenStudy (anonymous):

(2,-2) are Cartesian co-ordinates.

OpenStudy (anonymous):

yw